### v1

Abstract
Context:
Aims:A sample of SiO-masing late-type stars located in the inner Galaxy is analyzed with the goal of better constraining their obscuration.
Methods: This reference sample allows us to define mathematical relations between their dereddened infrared colours and the observed colours (e.g. K-[8], K-[24]).
Results:The derived equations define a property (a locus) of these late-type stars. Therefore, they enable us to derive the interstellar extinction. With estimated spectral types, it is possible to decompose the total extinction in the two components (interstellar and envelope extinction).
Conclusions:These relations are very useful to classify extremely obscured late-type stars located in the inner Galaxy. Estimating the two extinction components is performable on an individual late-type star, independently of its surrounding, and also when a few mid-infrared measurements are available.
Discourse on infrared colours and bolometric corrections of SiO masing stars in the inner Milky Way.
Key Words.:
circumstellar matter – Galaxy: stellar content – infrared: stars – stars: late-type – Methods: data analysis1 Introduction
SiO masing stars are usually of late-spectral types, later than M4-M5. Dusty circumstellar envelopes surround the central star, where maser emission from SiO, OH, and HO molecules may originate. The starlight is absorbed, reprocessed, and reemitted at infrared wavelengths by the dusty envelope. In other words, the envelope represents a deforming/transforming mirror of the original stellar energy distribution (SED). The effect is dependent on both the interstellar absorption caused by a diffuse medium along the line of sight and on the optical depth of the envelope, which can range from zero to an equivalent of about A= 100 mag. In the inner Galaxy, the interstellar extinction can likewise reach A= 30–40 mag. Due to the severe dust obscuration and the patchy interstellar extinction, the reddening, distances, and luminosities of late-type stars cannot be easily inferred.
To characterize the SED of an asymptotic giant star (AGB) star, a decomposition of its total extinction into the interstellar component and the circumstellar component is required. Messineo et al. (2005) measured the total extinction (A) for a sample of masing AGB stars and broke it down into A (interstellar) and A (envelope). Their method was straightforward and relied on multi-wavelength (single-epoch) photometric measurements of the targeted AGB star, covering the near- and mid-infrared window. They also assumed that the median extinction of nearby giants could be used to approximate the interstellar extinction.
Stellar colours and bolometric corrections (BC) of an AGB star vary depending on the properties of its envelope, e.g. extent, density, and mass loss rate. Here we present correlations between stellar colours, bolometric corrections, and envelope types obtained for the sample of SiO masing stars from Messineo et al. (2018). The advantage of this sample is that it covers a wide range of interstellar extinction and infrared stellar colours. We used it to elaborate on the intrinsic colours of cold stars in Messineo (2004), Messineo et al. (2005), and Messineo et al. (2012). This time we use this reference sample to develop a new technique. We establish correlations between the intrinsic infrared colours and the observed infrared colours and envelope types, assuming that we know the Galactic interstellar absorption curve. By solving a system of equations, it is, then, technically possible to decompose the extinction into the interstellar and envelope components, independently of the star surrounding. It is possible to model the BCs as a function of the intrinsic stellar colors and to determine the apparent bolometric magnitudes.
This idea is especially helpful in the innermost obscured regions of the Milky Way, where it is not possible to study the entire SED, and individual measurements in the mid-infrared are often of poor quality (e.gg., because of confusion). If one knows the nature of the star, for example, because of its maser emission, one can use the established equations for the K-[24] or K-[8] colours to determine the extinction, characterize the SED of the star, and estimate the apparent bolometric magnitude.
We describe the stellar sample and the available parameters in Sect. 2. Extinction-free colors are defined and the relationships between them and the dereddened colors are described in Sect. 3. The existence of these relations allows us to determine the interstellar extinction. Using an average spectral type (such as M6), one can also estimate the envelope extinction given the interstellar extinction.
Synthetic infrared colors derived from the DUSTY models are shown and compared with the observational data in Sect. 4.
The empirical estimates of extinction are compared with previously published estimates in Sect. 5.
2 The sample: The catalog of Messineo et al. (2018)
The sample consists of 572 mass-losing late-type stars which were searched for SiO maser emission (Messineo et al. 2002, 2018). The sources are located at low latitudes (). The sample had been created using near- and mid-infrared photometric measurements. The collected measurements cover from 0.9 m to 24 m and come from the publicly available surveys 2MASS, MSX, WISE, and GLIMPSE111 2MASS stands for Two Micron All Sky Survey (Skrutskie et al. 2006), DENIS for Deep Near Infrared Survey of the Southern Sky (Epchtein et al. 1994), MSX for Midcourse Space Experiment (Price et al. 2001), WISE for Wide-field Infrared Survey Explorer (Wright et al. 2010), and GLIMPSE for Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (Churchwell et al. 2009). (Messineo et al. 2018). Each photometric data point is from a single epoch. Several stellar parameters were estimated by Messineo et al. (2005) and Messineo et al. (2018).
Total A determination. The total extinction was estimated by Messineo et al. (2005, 2018), in the K versus K or K diagram, by shifting the SiO masing stars onto the 47 Tuc giant branch along the reddening vector – which implies by assuming a late M-type (M4-M5). The values of A span from 0 to 5.5 mag.
Interstellar A determination. The mean of surrounding stars is used as a proxy for A (Messineo et al. 2005, 2018). This assumption relies on the strong hypotheses that red giants and AGBs share the same Galactic distribution and that the target star is situated at the peak of the stellar density along the line of sight. Masing AGB stars are good tracers of the thin central bar, as shown by Habing et al. (2006). The thin disk population and the thick disk population do, in fact, cohabit in the disk; however, at low latitudes, the thick disk’s contribution is less than 10% (Lee et al. 2011).
Apparent bolometric magnitudes. Messineo et al. (2018) computed apparent bolometric magnitudes by directly integrating under the stellar SEDs from 0.9 m to 24 m. At longer wavelengths, the flux was linearly extrapolated to zero. At short wavelengths, the flux was estimated with a blackbody of 3000 K.
Infrared Bolometric corrections (BC). For a given photometric band, the BC represents the correction to be added to the dereddened apparent magnitude in order to estimate the stellar apparent bolometric magnitude (m), =m-m. There is a tight relation between the infrared BC values and the intrinsic stellar colours of Messineo et al. (2018). Second-order polynomial fits to the data are given in the Appendix.
2.1 Large amplitude variables
The sample of Messineo et al. (e.g. 2005) is made of Mira-like stars. Average photometric measurements should ideally be used to infer average colours and luminosities due to large periodic light variations. Amplitudes of Miras in the K band can reach up to 2 mag, while, at longer wavelengths, they are substantially smaller (e.g., at 15 m they are about 50% than those in K). Nevertheless, average quantities of the entire sample generate quantities typical of the actual average value (Messineo et al. 2004), based on the random phase of each star (absence of synchronous pulsations). Using DENIS, 2MASS, MSX, and WISE photometric flags, Messineo et al. (2018) found that at least 74% of the sample is made of long-period variables (LPV).
282 out of 572 stars have matches in the Gaia DR3 main catalog (Gaia Collaboration et al. 2023), 249 have a measured G-band magnitude; of those, 72% (179) are classified as variables; the G-band values range from 9 to 20 mag, there are from 10 to 62 G-band observations per star, from 2 to 60 BP-band observations, and from 7 to 65 RP-observations. So far, Gaia DR3 has released photometric data covering 34 months of observations (Eyer et al. 2023). Two scanning laws were used; the Ecliptic Pole Scanning Law (EPSL) during the first 28 days and the Nominal Scanning Law (NSL) afterward; the observational cadence depends on the period of the spin axis (6 h) and the precession around the Sun (63 d); the scanning law favours the Ecliptic poles. At the end of the six-year mission, each star is supposed to have an average of usable photometric transits.
For the targets detected as Gaia variables, the differences between the highest and lowest G-band measurements range from 0.12 mag to 3.54 mag (Eyer et al. 2023). Those marked as with “NOT_AVAILABLE” information on variability are fainter (mostly from G=18 to 21 mag) with a peak at G=20 mag, and have from 0 to 43 G-band observations, from 0 to 19 BP-band observations, and from 0 to 38 RP-observations. The differences between the highest and lowest G-band magnitudes range from 0.0 mag to 0.67 mag.
A fraction of 55% of Gaia matches are variables in Gaia and also in Messineo et al. (2018) based on infrared measurements. Gaia adds 16% of new variables (not detected with infrared flags).
The amplitude, or the difference between the minimum and maximum, is typically more than G=0.8 mag for Mira AGBs (e.g., Lebzelter et al. 2023; Messineo 2022). With variations greater than 0.8 mag, 74% of the Gaia variables are most likely Miras. But, only 86 of these 179 Gaia variables – whose amplitudes and periods range from 0.12 mag to 3.34 mag and their periods from 42 d to 808 d – are currently listed in the Gaia catalog of classified LPVs (Lebzelter et al. 2023). The Gaia catalog of LPVs includes only stars with a 5-95% quantile range larger than 0.1 mag in the G-band.
Since there is no discernible difference between the population of Gaia-identified LPVs and those not designated as variables for the purposes and diagrams displayed in this study, the fits and diagrams presented in this paper employ the whole sample of stars in Messineo et al. (2018).
3 Extinction
Messineo et al. (2012) have used
the interstellar extinction-free parameters and , where
, and
222The 2.69 coefficient
is obtained using the effective extinction values of Indebetouw et al. (2005).
to study the properties of different types of late-type stars.
The extinction-free colours are an observed quantity.
Messineo et al. (2012) show that the intrinsic MSX K colours,
(K, of SiO masing stars well correlate
with the parameter, as well as
with their BC() values.
The parameter spans a quite narrow range (1.5 mag), while
the parameter spans more than 10 mag (from 0 to 12 mag) and
is a very sensitive meter of circumstellar obscuration.
This correlation is of great advantage because it allows us to decompose the total A in the envelope component, A(env), and the interstellar component, A(int), and to obtain estimates of the stellar bolometric magnitudes. In this exercise, we revisit this idea to determine A(int) and A(env) quantitatively by taking the following steps:
Adopt an interstellar extinction curve.
Determine the Q values for stars in Messineo et al. (2005) that have A(int) values.
Verify the existence of mathematical relations between the Q values and, independently obtained, intrinsic colours.
For stars with unknown A(int), use these relations plus the observed Q values to estimate the intrinsic colours and A(int).
Using A and A(int), estimate the circumstellar extinction A(env).
Obtain constraints on the circumstellar extinction curve, as suggested by the kind referee.
3.1 Adopted interstellar extinction ratios
3.1.1 Near-infrared extinction power law
At near-infrared wavelength, the Galactic extinction curve is usually approximated with a power law, A and it appears the same in every direction. determinations in the literature range from 1.61 to 2.2. An index was measured by Rieke & Lebofsky (1985). A model with was favored in Messineo et al. (2005) based on an analysis of the 2MASS colors of late-type stars in the inner Galaxy. measurements larger than 2.0 were published in a number of works between 2006 and 2011 (e.g. Nishiyama et al. 2006; Stead & Hoare 2009). In the nuclear Disk, Fritz et al. (2011) reported a value of 2.1. For late-type stars at the Galactic center which suffer about 2 mag of reddening in band K, the increase of index from = 1.61 to = 1.9 implied a decrease of A of 0.7 mag. A change of index from = 1.9 to = 2.1 implies another decrease of about 0.35 mag. A higher decreases the extinction and makes the stars fainter.
In Messineo et al. (2005), an index of 1.9 was adopted, A/E(J-Ks)=0.537 and A/E(H-Ks)=1.493, and, ::=2.86:1.66:1.00. Using the same calculation performed in Messineo et al. (2005), for an index of 2.1, it is obtained that A/E(J-Ks)=0.455 and A/E(H-Ks)=1.311, and, ::=3.208:1.766:1.000. Within errors, the latter ratios are consistent with the empirical measurements of Wang & Chen (2019): ::=3.115:1.679:1.000. A/E(J-Ks)=0.473 and A/E(H-Ks)=1.472. At the Galactic center, Nogueras-Lara et al. (2020) measured = 1.87 0.03 and = 1.84 0.03, the power law with index= yields = 1.82 and = 1.77; while the Wang’s ratios yield = 1.85 and = 1.68. Here, a single power law with index= is adopted.
3.1.2 Mid-infrared extinction power law
The average effective extinction values for the filters of the MSX, WISE, and GLIMPSE surveys were estimated by convolving the extinction curves with a 3200 K giant of Allard et al. (2011) and the filter response curves.
As in Messineo et al. (2005), the parametric extinction curve of Rosenthal et al. (2000) (Curve 2), which was obtained in the OMC-1 cloud with H observations, was used along with two other curves. Curve 3 is a modified version of the Rosenthal curve in the 3-8 m region to incorporate the ratios from H-line observations in the Galactic center region of Fritz et al. (2011). Curve 4 was derived by Gordon et al. (2021) with Spitzer spectroscopic data of early-type stars located in the direction of a diffuse medium. The stars used by Gordon et al. are located at A 5 mag. The average effective extinction values from Curve 2, Curve 3, and Curve 4 are listed in Table 1.
The effective extinction ratios may vary at mid-infrared wavelengths. The peak of the silicate feature around 8-9 m is 1-1.3 in the Galactic center, but it is only 0.7 in the solar neighbouring stars observed by Spitzer. The Gordon curve can be regarded as a lower limit curve, while the Galactic center curve represents an upper limit curve.
The effective extinction ratios for the WISE and GLIMPSE filters were also determined by Wang & Chen (2019) using broad-band photometry of red clump stars, which are primary indicators of distances due to their known luminosity and colours. These are also shown in the Table 1 and seem to agree well with our predictions from the curves.
Filter | Curve 2 (Ros) | Curve 3 (Ros+Fritz) | Curve 4 (Gordon) | Wang | C (2) | C (3) | C (4) | ||
A/A | A/A | A/A | A/A | ||||||
[m ] | [m] | ||||||||
MSX A | 8.28 | 4.0 | 0.380 | 0.490 | 0.340 | 4.797 | 5.831 | 4.506 | |
MSX C | 12.1 | 2.1 | 0.491 | 0.491 | 0.320 | 5.843 | 5.843 | 4.374 | |
MSX D | 14.6 | 2.4 | 0.292 | 0.292 | 0.219 | 4.201 | 4.201 | 3.808 | |
MSX E | 21.3 | 6.9 | 0.405 | 0.405 | 0.300 | 4.998 | 4.998 | 4.249 | |
Glimpse [3.6] | 3.55 | 0.75 | 0.435 | 0.447 | 0.486 | 5.264 | 5.378 | 5.786 | |
Glimpse [4.5] | 4.49 | 1.01 | 0.298 | 0.402 | 0.351 | 4.237 | 4.973 | 4.582 | |
Glimpse [5.8] | 5.73 | 1.42 | 0.196 | 0.368 | 0.250 | 3.699 | 4.706 | 3.965 | |
Glimpse [8.0] | 7.87 | 2.93 | 0.295 | 0.408 | 0.266 | 4.218 | 5.024 | 4.052 | |
WISE W1 | 3.35 | 0.66 | 0.475 | 0.479 | 0.525 | 5.665 | 5.708 | 6.261 | |
WISE W2 | 4.60 | 1.04 | 0.284 | 0.398 | 0.337 | 4.154 | 4.940 | 4.486 | |
WISE W3 | 11.56 | 5.51 | 0.527 | 0.534 | 0.376 | 6.288 | 6.382 | 4.766 | |
WISE W4 | 22.09 | 4.10 | 0.394 | 0.394 | 0.294 | 4.908 | 4.908 | 4.213 |
WISE filter specifications are from Jarrett et al. (2011).
SPITZER filter specifications are from Fazio et al. (2004).
MSX filter wavelengths are from
() This column is based on the interstellar extinction curve obtained by Gordon et al. (2021) using stellar spectra taken with the Spitzer satellite.
() Values obtained with broad-band photometry by Wang & Chen (2019).
The C=C+C values are calculated using Eq. (2). C(2) refers to Curve 2, C(3) refers to Curve 3, and C(4) refers to Curve 4.
3.2 Definitions of the Q parameters
Based on the parameters and of Messineo et al. (2012), we can analogously calculate an interstellar extinction-free parameter Q for each filter and each interstellar extinction curve (Table 1):
where
and
where
By combining Eqs. 1 and 3, one obtains:
The constants can be estimated with an extinction curve. Curve 3 applies to highly obscured regions. The Gordon’s curve (for a diffuse interstellar gas) can be regarded as the Galactic lower limit (smaller peak at 10 m).
The filters GLIMPSE [8], WISE W3, and MSX A contain the 10 m silicate emission. The MSX D and WISE W4 filters appear to be particularly useful as they are not affected by the silicate emission, and show small variations of values in the three curves.
By definition, Eq. 5 yields an identical value when calculated with the observed magnitudes and the dereddened magnitudes (see Appendix). This allows us to estimate the intrinsic colours.
3.3 Q-intrinsic-colours curves for stars with known A(int)
Using the stars by Messineo et al. (2018) as a reference sample, a series of equations is set up to express the intrinsic colours as a function of the parameters.
The parameter Q correlates with the dereddened colour, e.g. (see Fig. 1). The fits are made with third-grade polynomial fits. It follows that
The intrinsic colours are calculated with the median A values of the surrounding stars (Messineo et al. 2005), which are rescaled to a power law with an index of , and the assumed effective extinction ratios. In Messineo et al. (2004) and Messineo et al. (2005), it is shown that the SiO masing stars are rare and the majority of the surrounding stars are normal giants. The large interval spanned by Q is due to the intrinsic colour range of the stars (stellar SEDs). These bright late-type stars are mostly mass-losing Mira AGBs with spectral types later than M4.
For each filter, the constants of the fits are listed in Table 2, when using Curve 3 of Messineo et al. (2005), and in Table 3 for the extinction curve of Gordon et al. (2021).
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Coef n0 | |
QA | (KsA)o | 0.18250.0319 | 0.00060.0028 | 0.00000.0001 | 0.16860.1108 | 0.15 |
QC | (KsC)o | 0.14530.0456 | 0.00210.0026 | 0.00000.0000 | 0.26360.2535 | 0.15 |
QD | (KsD)o | 0.22960.0261 | 0.00330.0010 | 0.18460.1619 | 0.20 | |
QE | (KsE)o | 0.18750.0162 | 0.00150.0004 | 0.21330.1474 | 0.18 | |
QW3 | (KsW3)o | 0.13550.0162 | 0.00260.0009 | 0.00040.0000 | 0.22000.0937 | 0.15 |
QW4 | (KsW4)o | 0.08870.0374 | 0.00810.0020 | 0.00010.0000 | 0.71270.2212 | 0.18 |
Q58 | (Ks-)o | 0.28530.0377 | 0.00130.0075 | 0.00000.0004 | 0.23010.0519 | 0.20 |
Q80 | (Ks-)o | 0.23510.0331 | 0.00180.0047 | 0.00000.0002 | 0.17360.0669 | 0.18 |
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Coef n0 | |
QA | (KsA)o | 0.26110.0501 | 0.00030.0067 | 0.00010.0003 | 0.31720.1107 | 0.20 |
QC | (KsC)o | 0.20570.0817 | 0.00470.0071 | 0.00000.0002 | 0.33860.2996 | 0.20 |
QD | (KsD)o | 0.26940.0318 | 0.00390.0015 | 0.25910.1696 | 0.22 | |
QE | (KsE)o | 0.23270.0228 | 0.00200.0008 | 0.27810.1660 | 0.21 | |
QW3 | (KsW3)o | 0.19420.0271 | 0.00520.0021 | 0.00010.0000 | 0.33580.1038 | 0.20 |
QW4 | (KsW4)o | 0.10060.0491 | 0.01260.0032 | 0.00030.0000 | 0.81900.2360 | 0.21 |
Q58 | (Ks-)o | 0.35990.0414 | 0.00230.0128 | 0.00000.0010 | 0.47670.0389 | 0.24 |
Q80 | (Ks-)o | 0.30600.0416 | 0.00520.0093 | 0.00040.0006 | 0.39770.0538 | 0.23 |
3.4 A and (K) colours derived with Q and the Q-intrinsic-colours curves.
The Q-intrinsic-colours curves derived in the above section can be used to obtain estimates of the (K) colours for stars with unknown parameters.
The following formula can be used to estimate the interstellar A as a function of the observed and dereddened colors once the intrinsic colors have been identified with Q:
The term results from Eq. 6 and the coefficients from Table 2 or 3 by entering the Q. The term is the observed colour. The extinction ratios are listed in Table 1.
The Curve 3 of Messineo et al. (2005) was used first. The four independent A obtained with the MSX magnitudes were averaged; 0.05 mag. We repeated the calculation with the WISE dataset and the GLIMPSE dataset. The A values from the W3 and W4 WISE bands give =0.03 mag, while the A values from GLIMPSE [5.8] and [8.0] data yield =0.03 mag. There are 135 stars with good photometry and estimates of A with MSX and WISE datasets; the average difference is 0.04 mag and the is 0.20 mag, as shown in Fig. 2. The 111 A estimates of good photometric quality in common between the GLIMPSE and WISE estimates yield =0.15 mag.
Identical A values are measured (within errors) after repeating the full computation using Gordon’s effective ratios of Table 1. The same A values are obtained as long as the dereddened colours and the are performed consistently with the same interstellar curve.
We measured A by analyzing the correlations between Q and the intrinsic colours. The A values of individual stars are obtained independently of the stellar environment, anywhere in the Galactic Disk.
In Fig. 3, we compare the original estimate of A (via the surrounding stars) and the one obtained with the fits and Q. Using the MSX data, the average A(fit)-A(surrounding)=0.14 mag with a mag and a median of 0.07 mag; using the WISE data the average A(fit)-A(surrounding)=0.04 mag with a mag and a median of 0.00 mag; using the GLIMPSE data A(fit)-A(surrounding)=0.05 mag with a mag and a median of 0.03 mag. This method allows for an accuracy of 0.35 mag, which corresponds to the scatter of the data points in the diagram K versus K of Fig. 1 (the near- and mid-infrared measurements are not acquired simultaneously).
3.5 Estimates of A as A-A
The K colours of naked giants were calculated by synthetic photometry of theoretical spectra. For this purpose, the publicly available NextGen spectra of Allard et al. (2011) with T=3500, 3200, 2600 K, log(g)=1, and solar metallicity were used333The spectra are distributed by the Virtual Observatory SED analyzer (Bayo et al. 2008). The bt-nextgen_agss2009 (gas only) were retrieved.. The synthetic spectra were generated with the PHOENIX atmospheric code in a spherical symmetry. The synthetic spectra reproduce well the TiO and VO molecules, which dominate the optical spectrum, and water vapour and CO molecules in the infrared spectrum. The molecular list was refined to reproduce the atmospheric lines of brown dwarfs well (Allard et al. 2011; Allard 2014). The infrared colours are obtained by convolving the spectrum with the filter and calculating the mean flux density. The adopted zero points come from Cohen et al. (2003, 2MASS), Egan et al. (2003, MSX), Jarrett et al. (2011, WISE), and Fazio et al. (2004, GLIMPSE). The synthetic infrared colours are listed in Table 4.
The average differences between the A(tot) calculated with the synthetic (K) colours and those obtained with the colours of Koornneef (1983) are +0.05 mag with =0.00 mag (M4) and +0.14 mag with =0.00 mag (M6). While the average differences between the A(tot) calculated with the synthetic (K) colours for a 3500 K (M4) and a 3200 K (M5-6) star are +0.015 mag with =0.006 mag.
The A(int) values were re-determined using the fiducial colour sequences (Q, as described in Sec. 3.
Using the derived A and A, we could estimate the envelope extinction A for 324 targets. A=AA ranges from 0 to 2.5 mag, as shown in Fig. 4. The envelope optical depth at 0.55 m, , can be estimated as A and ranges from 0 to 32.
T | [K] | 3500 | 3200 | 2600 |
Spectral type | M4 | M5-6 | M9 | |
[mag] | 1.198 | 1.245 | 0.967 | |
[mag] | 0.231 | 0.244 | 0.222 | |
[mag] | 0.000 | 0.000 | 0.000 | |
[mag] | 0.115 | 0.104 | 0.320 | |
[mag] | 0.038 | 0.449 | ||
[mag] | 0.032 | 0.051 | 0.473 | |
[mag] | 0.194 | 0.161 | 0.596 | |
[mag] | 0.052 | 0.041 | 0.504 | |
[mag] | 0.208 | 0.228 | 0.796 | |
[mag] | 0.216 | 0.242 | 0.822 | |
[mag] | 0.225 | 0.275 | 0.888 | |
[mag] | 0.081 | 0.009 | 0.151 | |
[mag] | 0.006 | 0.425 | ||
[mag] | 0.207 | 0.219 | 0.759 | |
[mag] | 0.218 | 0.271 | 0.887 |
3.6 Empirical circumstellar excess colours.
Excess ratios | mean_1 | sigma_1 | mean_2 | sigma_2 |
(Curve3) | (Curve3) | (Gordon) | (Gordon) | |
0.73 | 0.10 | 0.70 | 0.11 | |
1.02 | 0.05 | 0.97 | 0.04 | |
1.20 | 0.10 | 1.18 | 0.10 | |
0.24 | 0.22 | 0.31 | 0.23 | |
0.54 | 0.46 | 0.58 | 0.54 | |
1.44 | 0.39 | 1.50 | 0.47 | |
0.08 | 0.67 | 0.23 | 0.58 | |
0.28 | 0.89 | 0.31 | 0.91 | |
1.24 | 0.25 | 1.20 | 0.30 |
The mean_1 and sigma_1 values are obtained after having dereddened the observed colours with the A values and the interstellar extinction Curve 3 of Messineo et al. (2005).
Excess ratios | mean_1 | sigma_1 |
(Curve3) | (Curve3) | |
0.68 | 0.43 | |
1.04 | 0.17 | |
1.26 | 0.46 | |
0.34 | 0.78 | |
0.72 | 0.53 | |
3.19 | 36.22 | |
0.05 | 2.97 | |
0.30 | 2.91 | |
1.30 | 0.80 |
The mean_1 and sigma_1 values are obtained after having dereddened the observed colours with the A values and the interstellar extinction Curve 3 of Messineo et al. (2005).
As suggested by the kind referee, the naked colours listed in Table 4 can be useful to determine the colour excess ratios and to add constraints on the circumstellar extinction curve.
A circumstellar colour excess is defined as
where is the interstellar-dereddened colour and is the expected colour for a star of similar parameters (e.g. gravity, metallicity, and temperature) but without an envelope. For the values, the synthetic colours are taken from a solar model at 3200 K (Table 4).
In Fig. 5, the E(Ks-[D]) colour excess is plotted against the E(K-[A]), E(Ks-[C]), and E(Ks-[E]) colour excess. The Figs. 6 and 7 show similar diagrams for the WISE and GLIMPSE filters. There is a tight correlation between the MSX excess colours, as well as between the E(Ks-[W4]) and E(Ks-[W3]) excess colours, and the E(Ks-[8.0]) and E(Ks-[5.8]) excess colours. On the contrary, a large scatter appears in the diagrams showing the E(Ks-[W1]), E(Ks-[W2]), E(Ks-[3.6]), and E(Ks-[4.5]) excess colours in Figs. 6 and 7. Indeed, the stellar spectrum of cold stars short-ward of 4 m is dominated by strong molecular bands (Bedijn 1987). Furthermore, the data used here are from a single-epoch, and near-infrared data and mid-infrared were not taken simultaneously. The radial pulsation expands the atmosphere, and creates an extra layer of molecules (CO, HO, CO, and SiO) above the atmosphere. The region from 2.4 to 4 m is dominated by strong water vapour, which is the main source of opacity. The water vapour absorption changes dramatically with the pulsation phase (Matsuura et al. 2002). This extended layer of molecules was analyzed in detail in the work of Matsuura et al. (2002) and Cami (2002). The larger scatter of data points shown in Fig. 6 and 7, with the filters W1 and W2, and with the filters [3.6] and [4.5], indicates the presence of strong absorption by water.
A tighter correlation of the colour excess E(K-D) with E(K-A), E(K-C), and E(K-E), confirms that the filters A, C, D, and E are dominated by dust emission, as well as W3, W4, [5.8], and [8.0]. As already written by Volk & Kwok (1988) for the LRS spectra, “in the region from 8-23 m the emergent spectrum will directly reflect the shape of the opacity function”.
Using Eq. (8), the excess ratios can also be calculated. The median excess ratios and standard deviations are tabulated in Table 5. The excess ratios in Tables 5 and 6 which were calculated with the MSX bands, WISE W3 and W4, and GLIMPSE [8.0] bands, appear well described by constant values and can provide constraints about the circumstellar extinction curve of the sampled O-rich AGB stars (e.g., Bedijn 1987; Volk & Kwok 1988; Ossenkopf et al. 1992).
4 A grid of black body models with mass loss
We assume a fixed temperature of 3200 K (M6), the astronomical silicate grains of Ossenkopf et al. (1992), and use the DUSTY code of Ivezic et al. (1999) to generate a set of SED models with optical depth, , from 0 to 40. The model consists of a spherical shell with density decreasing with . The dust condensation radius was set to 1000 K.
In the diagram versus K, the generated model describes the data points of Messineo et al. (2018) well, as shown in Fig. 8. In the diagram versus K, the models display smaller than the data points. Indeed, the W3 filter encompasses the 10 m feature and also the 13 m dust bump; these features are not well reproduced by the silicate of Ossenkopf et al. (1992). Indeed, a much better agreement with the colour is found using a mixture of cold and warm silicates and amorphous alumina (Suh 2021). Fig. 8 shows how changing the results in a a colour-colour track, given a stellar temperature and a dust type. Each point on the track represents a specific tau-point. However, when varying the dust size, the scale changes. This can be seen in Fig. 9, where the dereddened (K is plotted against the envelope. A larger maximum size results in a larger for a given colour. In recent literature, the standard distribution of grain sizes, MNR, is usually adopted with maximum sizes from 0.1 to 1. m (e.g. Olofsson et al. 2022; van Loon et al. 2005; Wang et al. 2021) or fixed sizes of 0.1-0.5 m (e.g. Blommaert et al. 2018; Suh 2021).
Circumstellar effective extinction ratios are highly dependent on the maximum grain size. We estimated that using the DUSTY code, the astronomical silicate grains by Ossenkopf et al. (1992), and an MNR distribution from 0.005 to 0.75 m. For a maximum size of 0.45 m, , and when the maximum size is 0.25 m. Three circumstellar extinction curves with the MNR distribution and maximum sizes of 0.75, 0.45, and 0.25 m are plotted in Fig. 10 to illustrate this. The curves show different peak values of the 9.7 feature (1.4, 0.7, and 0.4, respectively) when normalizing at 2.12 m. Compared to a curve with a maximum grain size of 0.25 m, an extinction curve with a maximum grain size of 0.75 m appears to have a lower 9.7 m peak opacity. The same curve appears steeper in the near-infrared when normaled at 9.7 m, than the curve with a maximum grain size of 0.25 m.
In Table 7, theoretical excess ratios are calculated by convolving the curves with a maximum grain size of 0.75 m and 0.25 m with the filter profiles and a cool spectrum by Allard et al. (2011).
A comparison of the theoretical excess ratios of Table 7 with those observationally measured (Table 5), suggests that an MNR distribution with a maximum grain size of 0.25 m produced excess ratios that were smaller than those observed. An MNR distribution with a maximum grain size of 0.75 m provides a better agreement with the data.
Excess ratios | Temp | mean_1 | mean_2 | mean_3 | mean_4 | mean_4 |
[K] | (max_25) | (max_75) | (Ros_0.4) | (Ros_1.4) | (Ros_0.2) | |
3200 | 0.966 | 0.964 | 0.882 | 0.889 | 0.905 | |
3200 | 0.620 | 0.928 | 0.906 | 0.548 | 0.956 | |
3200 | 0.563 | 0.924 | 0.967 | 0.715 | 1.000 | |
3200 | 1.822 | 0.768 | 0.693 | 1.738 | 0.608 | |
3200 | 2.537 | 1.081 | 0.943 | 2.368 | 0.828 | |
3200 | 1.098 | 1.032 | 1.121 | 1.467 | 1.078 | |
3200 | 0.741 | 0.731 | 0.701 | 0.701 | 0.701 | |
3200 | 0.919 | 0.916 | 0.872 | 0.872 | 0.872 | |
3200 | 0.791 | 0.971 | 1.006 | 0.806 | 1.060 | |
2600 | 0.935 | 0.956 | 0.875 | 0.858 | 0.903 | |
2600 | 0.617 | 0.928 | 0.905 | 0.539 | 0.956 | |
2600 | 0.557 | 0.922 | 0.968 | 0.709 | 1.001 | |
2600 | 1.886 | 0.782 | 0.695 | 1.798 | 0.608 | |
2600 | 2.591 | 1.084 | 0.944 | 2.442 | 0.826 | |
2600 | 1.107 | 1.034 | 1.123 | 1.490 | 1.078 | |
2600 | 0.748 | 0.738 | 0.703 | 0.703 | 0.703 | |
2600 | 0.922 | 0.917 | 0.873 | 0.873 | 0.873 | |
2600 | 0.780 | 0.968 | 1.004 | 0.793 | 1.060 |
Circumstellar excess ratios were calculated with the MSX, WISE, and GLIMPSE filter response curves, with a stellar spectrum of 3200 K (Allard et al. 2011), and extinction curves. The first column (max_25) refers to the extinction curve obtained with DUSTY adopting a maximum size of 0.25 m (see text), which is similar to Rosenthal’s curve normalized at 2.12 m and with a silicate peak at 1.4. The second column (max_75) refers to the extinction curve obtained with DUSTY adopting a maximum size of 0.75 m, which is close to Rosenthal’s curve normalized at 2.12 m and with a silicate peak at 0.4. The spectrum of a naked star of 3200 K is used to obtain the naked star colours.
In order to reproduce the observed colours, small values of the maximum grain sizes (0.1-0.25 m) are frequently reported in the literature. For instance, Volk & Kwok (1988) estimate for their sample of IRAS sources using dust grain sizes smaller than 0.25 m. However, as David & Papoular (1990) points out, observational data (IRAS data) only constrains their opacity curve in the mid-infrared region; for the near-infrared curve, the theoretical model by Draine (1985) is used. Bedijn (1987) uses a power law with an index for the opacity curve below 5 m, yielding .
Recently, Maercker et al. (2022) preferred larger grain sizes (2 m) to fit the far-infrared fluxes of Carbon AGB stars. Large dust grains with an average size of 0.5 m were detected in the envelopes of VY Canis Majoris (Scicluna et al. 2015) and of W Hydrae (Ohnaka et al. 2016) with optical polarimetric imaging.
5 Discussion
Relations between the de-reddened colours and the extinction-free colours Q, Q,Q,Q,..Q were constructed using a set of O-rich Mira-like stars with known interstellar extinction (Messineo et al. 2005). Such equations allow us to determine the interstellar extinction for every Mira-like star that has accessible near- and mid-infrared colors.
In order to verify the applicability and reliability of the new methodology in determining interstellar extinction, an analysis of existing catalogs of Miras with available and mid-infrared measurements is made.
The most obscured Miras often exhibit OH maser emission and are therefore called OH/IR stars.
Typically, their flux distribution is modeled with radiative transfer codes, and their interstellar extinction is derived from infrared extinction maps. For example, Olofsson et al. (2022) analyzed a sample of 22 AGB stars with OH masers near the Galactic center by using CO line observations taken with The Atacama large (sub-)millimeter array (Alma) array. The map of extinction by Gonzalez et al. (2018) and the extinction law by Nishiyama et al. (2009) were used. Archival photometry (from the 2MASS, VVV, and UKIDSS surveys 444VVV stands for the VISTA Variables in the Via Lactea survey (Soto et al. 2013). The UKIDSS Galactic Plane Survey (GPS) is described in Lucas et al. (2008)., the catalogues of Nogueras-Lara et al. (2019) and Nogueras-Lara et al. (2021)) and mid-infrared photometry from GLIMPSE surveys (5.8 and 8.0 um) could be retrieved from the VIZIER database for 8 stars of the Olofsson’s sample. The equations of Table 2 were used to make interstellar extinction estimates (A(int) from 0.90 to 2.0 mag). The two sets of A(int) linearly correlate with a standard deviation =0.49 mag and a mean difference of 0.09 mag (or 0.24 mag when rescaling the extinction to a power law with index = 2.1). The non-coevality of the near- and mid-infrared photometry accounts for the large ; the OH/IR stars have typical K amplitudes ranging from 0.9 to 2.5 mag (Messineo et al. 2020).
Using data from the ISOGAL survey (Omont et al. 2003), Ojha et al. (2003) determined interstellar extinction, mass-loss rate, and luminosity values for 321 late-type stars located in Bulge fields at (, ). The majority of the sampled stars are AGB stars because they are brighter than the stars at the tip of the red giant branch. The authors used DENIS and ISOGAL photometry at 7 and 15 m and the silicate models by Jeong et al. (2003). The new method is used to estimate A(int) for 228 stars in this sample based on GLIMPSE and 2MASS data. The Ojha’s A values are converted to A using a factor of 0.089 (Ojha et al. 2003), and, then, compared with the A obtained here. The mean difference is 0.09 mag and a mag.
Matsunaga et al. (2009) estimate the A(int) values of 52 Mira stars at the Galactic center using average near-infrared measurements and the extinction law by Nishiyama et al. (2006).
The average measurements along with the GLIMPSE [5.8] and [8.0] magnitudes are used to compute the A(int) values for 30 stars using the new method. The A(HK) from Matsunaga were converted to a power law with index = (by multiplying for 0.91) and range from 1.81 to 2.59 mag. An average difference of 0.1 mag is obtained and mag.
The established equations between intrinsic infrared colours and extinction-free colours (Q), which are listed in Table 2, are useful to automatically determine stellar obscuration. This set of equations can be used to select late-type stars in a given sky region, and to improve their estimates of luminosity.
Assuming a reasonable estimate of the total, the envelope extinction can be calculated as the difference between the total and the interstellar extinction. Mathematically, it is a straightforward calculation. In reality, estimating the circumstellar extinction is actually difficult because accurate spectral types are lacking, naked-star colors are unknown and affected by strong variable molecular absorption.
Future photometric multi-wavelengths and multi-epochs surveys and spectroscopic surveys, such as LSST, Gaia DR4, 4MOST, GALAH, will enable us to obtain spectral types and temperatures, and their variations. This will allow us to perform more accurate extinction calculations.
Appendix A BC and intrinsic colours versus Q.
In Messineo (2004), BC values as a function of infrared colours were obtained for the MSX filters. M values and dereddened WISE and GLIMPSE magnitudes from the revised catalog of Messineo et al. (2018) are here used to extend the BC computation to the WISE and Glimpse filters. Polynomial fits are listed in Table 8.
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Y-axis Y-Fit | X-range | |
(mag) | (mag) | ||||||
(KsA)o | BC(Ks) | 2.263 0.039 | 0.691 0.028 | -0.156 0.005 | -0.000 | 0.210 | |
(KsA)o | BC(A) | 2.263 0.039 | 1.691 0.028 | -0.156 0.005 | 0.000 | 0.210 | |
(KsC)o | BC(Ks) | 1.594 0.069 | 0.882 0.036 | -0.136 0.004 | -0.000 | 0.224 | |
(KsC)o | BC(C) | 1.594 0.069 | 1.882 0.036 | -0.136 0.004 | -0.000 | 0.224 | |
(KsD)o | BC(Ks) | 1.618 0.068 | 0.861 0.036 | -0.131 0.004 | 0.000 | 0.223 | |
(KsD)o | BC(D) | 1.618 0.068 | 1.861 0.036 | -0.131 0.004 | 0.000 | 0.223 | |
(KsE)o | BC(Ks) | 1.391 0.131 | 0.852 0.057 | -0.111 0.006 | 0.000 | 0.230 | |
(KsE)o | BC(E) | 1.391 0.131 | 1.852 0.057 | -0.111 0.006 | 0.000 | 0.230 | |
(KsW1)o | BC(Ks) | 2.943 0.028 | 0.172 0.042 | -0.245 0.013 | 0.000 | 0.285 | |
(KsW1)o | BC(W1) | 2.943 0.028 | 1.172 0.042 | -0.245 0.013 | 0.000 | 0.285 | |
(KsW2)o | BC(Ks) | 2.839 0.032 | 0.264 0.028 | -0.136 0.005 | 0.000 | 0.270 | |
(KsW2)o | BC(W2) | 2.839 0.032 | 1.264 0.028 | -0.136 0.005 | -0.000 | 0.270 | |
(KsW3)o | BC(Ks) | 2.113 0.057 | 0.659 0.032 | -0.121 0.004 | -0.000 | 0.255 | |
(KsW3)o | BC(W3) | 2.113 0.057 | 1.659 0.032 | -0.121 0.004 | -0.000 | 0.255 | |
(KsW4)o | BC(Ks) | 1.490 0.084 | 0.799 0.037 | -0.104 0.004 | 0.000 | 0.271 | |
(KsW4)o | BC(W4) | 1.490 0.084 | 1.799 0.037 | -0.104 0.004 | -0.000 | 0.271 | |
(Ks)o | BC(Ks) | 3.094 0.021 | -0.064 0.024 | -0.153 0.011 | 0.000 | 0.316 | |
(Ks)o | BC( | 3.094 0.021 | 0.936 0.024 | -0.153 0.011 | -0.000 | 0.316 | |
(Ks)o | BC(Ks) | 2.986 0.024 | 0.255 0.037 | -0.180 0.012 | -0.000 | 0.293 | |
(Ks)o | BC() | 2.986 0.024 | 1.255 0.037 | -0.180 0.012 | -0.000 | 0.293 | |
(Ks)o | BC(Ks) | 2.637 0.034 | 0.559 0.034 | -0.169 0.007 | -0.000 | 0.244 | |
(Ks)o | BC() | 2.637 0.034 | 1.559 0.034 | -0.169 0.007 | -0.000 | 0.244 | |
(Ks)o | BC(Ks) | 2.490 0.044 | 0.602 0.036 | -0.146 0.007 | -0.000 | 0.249 | |
(Ks)o | BC() | 2.490 0.044 | 1.602 0.036 | -0.146 0.007 | 0.000 | 0.249 |
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Y-axis Y-Fit | X-range | |
(mag) | (mag) | ||||||
(KsA)o | BC(Ks) | 2.540 0.028 | 0.632 0.021 | 0.156 0.003 | 0.000 | 0.158 | |
(KsA)o | BC(A) | 2.540 0.028 | 1.632 0.021 | 0.156 0.003 | 0.000 | 0.158 | |
(KsC)o | BC(Ks) | 1.869 0.051 | 0.857 0.027 | 0.140 0.003 | 0.000 | 0.169 | |
(KsC)o | BC(C) | 1.869 0.051 | 1.857 0.027 | 0.140 0.003 | 0.000 | 0.169 | |
(KsD)o | BC(Ks) | 1.834 0.055 | 0.846 0.028 | 0.132 0.003 | 0.000 | 0.176 | |
(KsD)o | BC(D) | 1.834 0.055 | 1.846 0.028 | 0.132 0.003 | 0.000 | 0.176 | |
(KsE)o | BC(Ks) | 1.605 0.104 | 0.850 0.045 | 0.114 0.005 | 0.000 | 0.183 | |
(KsE)o | BC(E) | 1.605 0.104 | 1.850 0.045 | 0.114 0.005 | 0.000 | 0.183 | |
(KsW1)o | BC(Ks) | 3.088 0.026 | 0.154 0.037 | 0.239 0.011 | 0.000 | 0.248 | |
(KsW1)o | BC(W1) | 3.088 0.026 | 1.154 0.037 | 0.239 0.011 | 0.000 | 0.248 | |
(KsW2)o | BC(Ks) | 3.074 0.028 | 0.165 0.024 | 0.126 0.005 | 0.000 | 0.229 | |
(KsW2)o | BC(W2) | 3.074 0.028 | 1.165 0.024 | 0.126 0.005 | 0.000 | 0.229 | |
(KsW3)o | BC(Ks) | 2.459 0.047 | 0.562 0.027 | 0.117 0.004 | 0.000 | 0.218 | |
(KsW3)o | BC(W3) | 2.459 0.047 | 1.562 0.027 | 0.117 0.004 | 0.000 | 0.218 | |
(KsW4)o | BC(Ks) | 1.789 0.074 | 0.751 0.032 | 0.103 0.003 | 0.000 | 0.238 | |
(KsW4)o | BC(W4) | 1.789 0.074 | 1.751 0.032 | 0.103 0.003 | 0.000 | 0.238 | |
(Ks)o | BC(Ks) | 3.270 0.019 | 0.0790.022 | 0.162 0.009 | 0.000 | 0.265 | |
(Ks)o | BC( | 3.270 0.019 | 0.921 0.022 | 0.162 0.009 | 0.000 | 0.265 | |
(Ks)o | BC(Ks) | 3.158 0.021 | 0.199 0.032 | 0.177 0.010 | 0.000 | 0.243 | |
(Ks)o | BC() | 3.158 0.021 | 1.199 0.032 | 0.177 0.010 | 0.000 | 0.243 | |
(Ks)o | BC(Ks) | 2.884 0.027 | 0.462 0.027 | 0.162 0.006 | 0.000 | 0.198 | |
(Ks)o | BC() | 2.884 0.027 | 1.462 0.027 | 0.162 0.006 | 0.000 | 0.198 | |
(Ks)o | BC(Ks) | 2.781 0.034 | 0.506 0.028 | 0.141 0.005 | 0.000 | 0.201 | |
(Ks)o | BC() | 2.781 0.034 | 1.506 0.028 | 0.141 0.005 | 0.000 | 0.201 |
The BC values and the intrinsic colours are plotted versus the interstellar-extinction-free parameter Qin Figs. 11, 12, and 13 for the MSX, WISE, and GLIMPSE filters.
Appendix B and the jolly equation
This is the definition of (Ks-) colour excess:
By combining Eq. B2. and B.3, it is obtained
and, therefore,
where
The B.5 is the definition of the parameter.
The first member contains observed quantities,
while the second term only contains intrinsic colours.
This equivalence proves that does not depend
on interstellar extinction. It is an intrinsic colour of the star.
Analogously, a function can be defined with the colour.